# Asymptotic directions in the moduli space of curves

### v1

@@ 5. Deformations of rank $d<Cliff(C)$
Let $C$ be a smooth algebraic curve of genus $g$ and $\zeta\in H^{1}(T_{C})$ a deformation of rank $d$. Suppose one of the following assumptions is satisfied:
@@
So, we have $h^{0}(K_{C}\otimes A^{\vee})\geq 1$. We claim that $h^{0}(A)\geq 2$. In fact, if $h^{0}(A)\leq 1$, we would have $h^{0}(K_{C}\otimes A^{\vee})\geq h^{0}(E)-1=g-d$ and by Riemann Roch, we would get $deg(A)\leq d<\frac{g-1}{2}$, a contradiction. So $h^{0}(A)\geq 2$, and if $h^{0}(K_{C}\otimes A^{\vee})\geq 2$, $A$ contributes to the Clifford index and we have:
@@
as the locus of deformations of rank $d$. When $d=1$ it is the bicanonical curve.
@@
Let $C$ be a smooth curve of genus $g$, take an integer $d<Cliff(C)$ and an effective divisor $D$ of degree $d$. Then $K_{C}(-D)$ is projectively normal, so the multiplication map
@@
Notice that $h^{0}(D)=1$ since $Cliff(C)>d$. Moreover $K_{C}(-D)$ is very ample by Riemann Roch and by the assumption $Cliff(C)>d$. So by [15, Theorem 1], $K_{C}(-D)$ is projectively normal. ∎
@@
Let $C$ be a smooth curve of genus $g$, take an integer $d<Cliff(C)$, an effective divisor $D=\sum_{i=1}^{k}m_{i}p_{i}$ of degree $d$. Then any linear combination of all Schiffer variations $\xi^{n_{i}}_{p_{i}}$, with $n_{i}\leq m_{i}$ is not an asymptotic direction.
@@
We will prove the statement by induction on $d$. The case $d=1$ is already known (see Lemma 8.2).
By induction assume that no linear combination of all $\{\xi^{l_{i}}_{p_{i}}\ |\ l_{i}\leq m_{i},\ i<k,\ l_{k}\leq m_{k}-1\}_{i=1,...,k}$ is asymptotic.
@@
Let $C$ be a smooth curve of genus $g$, take an integer $d<Cliff(C)$. Then there are no asymptotic directions $\zeta\in H^{1}(T_{C})$ of rank $d$.
@@
    all tangent directions $\zeta$ of rank $d<\lfloor\frac{g-1}{2}\rfloor$ are linear combinations of Schiffer variations supported on an effective divisor of degree $d$.
@@
    There are no asymptotic directions of rank $d<\lfloor\frac{g-1}{2}\rfloor$.
@@
For the general curve of genus $g$ we characterise the image of the natural map ${\mathbb{P}}T_{C_{d}}\to{\mathbb{P}}(H^{1}(T_{C}))$ as the locus of deformations of rank $d$, for all $d<\lfloor\frac{g-1}{2}\rfloor$.
@@ 6. Double-split deformations
A deformation $\zeta\in H^{1}(T_{C})$ is said to be double-split if it is split with $\omega=s_{1}\tau$, $s_{1}\in H^{0}(L)$, $\tau\in H^{0}(K_{C}\otimes L^{\vee})$ and $\tau=s_{2}t$ with $s_{2}\in H^{0}(L)$ and $t\in H^{0}(K_{C}\otimes L^{-2})$.
@@
With the above notation, if $\zeta$ is double-split, we have
$II(Q)(\zeta\odot\zeta)=deg(L)\neq 0.$
@@
$II(Q)(\zeta\odot\zeta)=-\sum_{p\in Supp(D)}Res_{p}(g_{1}dg_{2}-g_{3}dg_{4})=-\sum_{p\in Supp(D)}Res_{p}(g_{1}dg_{2})=$
$=-\sum_{p\in Supp(D)}Res_{p}(xd(\frac{1}{x}))=\sum_{p\in Supp(D)}Res_{p}(\frac{dx}{x})=deg(D)=deg(L).$
@@ 7. Rank $d=Cliff(C)$ deformations
Taking gobal sections in (1), if we $\zeta$ has rank $d$, we get $h^{0}(E)=g-d+1$.
@@
Recall that if $\zeta\in H^{1}(T_{C})$ is a deformation of rank $d=Cliff(C)<\frac{g-1}{2}$ and $E$ is not globally generated, then by Theorem 5.1, $\zeta$ is a linear combination of Schiffer variations supported on a degree $d$ effective divisor.
@@
Assume $\zeta\in H^{1}(T_{C})$ is an infinitesimal deformation of rank $d$ on a curve $C$ of Clifford index $d\neq\frac{g-1}{2}$ and $g\geq 5$. Assume moreover that $Cliff(C)=Cliff(L)=gon(C)-2$, where $L$ is a $g^{1}_{d+2}$ on $C$. Assume that $\zeta$ is not a linear combination of Schiffer variations supported on a degree $d$ effective divisor.
@@
By the assumptions and by Remark 7.1, we know that $E$ is globally generated. Take to generic points $p,q\in C$, then since $g\geq 5$, there exists a nontrivial holomorphic section $\sigma\in H^{0}(E(-p-q))$.
@@
where $E^{\prime}=E\otimes{\mathcal{O}}_{C}(-p-q)$, so $h^{0}(E^{\prime})\geq g-d-3>0$. By the theorem of Segre-Nagata and Ghione, ([19] p. 84), there exists a subline bundle $A$ of $E^{\prime}$ such that $deg(A)\geq\frac{g-5}{2}$. We claim that $h^{0}(A)\geq 1$. In fact, if $h^{0}(A)=0$, by Riemann Roch we would have $h^{0}(K_{C}\otimes A^{\vee})=g-1-deg(A)\leq\frac{g+3}{2}$. So, since $p,q$ are generic points, $h^{0}(K_{C}\otimes A^{\vee}(-2p-2q))\leq\frac{g-5}{2}$. On the other hand, up to saturation we have the exact sequence
@@
We have $h^{0}(A(p+q))\geq 1$ and $deg(A(p+q))\geq\frac{g-1}{2}$. If $h^{0}(A(p+q))=1$, then $h^{0}(K_{C}\otimes A^{\vee}(-p-q))=g-2-deg(A)\leq\frac{g+1}{2}$, but $h^{0}(K_{C}\otimes A^{\vee}(-p-q))\geq h^{0}(E)-1\geq g-d$, so we get $d\geq\frac{g-1}{2}$, a contradiction. So $h^{0}(A(p+q))\geq 2$ and $h^{0}(K_{C}\otimes A^{\vee}(-p-q))\geq 2$, since $E$ is globally generated. So $A(p+q)$ contributes to the Clifford index and we have:
@@
hence $d\leq Cliff(A(p+q))\leq d$, and either $A(p+q)=L$, or $A(p+q)=K_{C}\otimes L^{\vee}$, where $L$ is any line bundle as in the statement. But $A(p+q)\neq L$, since $p,q$ are general and $A$ is effective. So $A(p+q)=K_{C}\otimes L^{\vee}$, and the vertical exact sequence induces an exact sequence on global sections. So it corresponds to a class of an extension $\epsilon\in H^{1}(K_{C}\otimes L^{-2})$ that induces the zero map $H^{0}(L)\to H^{1}(K_{C}\otimes L^{\vee})$. If $h^{0}(L^{\otimes 2})=3$, the multiplication map $Sym^{2}H^{0}(L)\to H^{0}(L^{\otimes 2})$ is surjective, so $\epsilon$ must be zero. So $E=(K_{C}\otimes L^{\vee})\oplus L$ and $\zeta$ is split.
@@
Assume now $h^{0}(3L)=4$, then the multiplication map
@@ 8. Rank one deformations
Assume $C$ is not hyperelliptic, $\zeta\in H^{1}(T_{C})$ a rank one deformation. If $E$ is not globally generated, then $\zeta=\xi_{p}$ is a Schiffer at a point $p$. If $C$ non trigonal, then $\forall p\in C$, $\xi_{p}$ is not an asymptotic direction. If $C$ is trigonal and $g\geq 6$, then $\xi_{p}$ is an asymptotic direction if and only if $p$ is a ramification point of the $g^{1}_{3}$.
@@
Assume $E$ is not globally generated, we conclude by Theorem 5.1 that $\zeta=\xi_{p}$ is a Schiffer at a point $p$.
If $g\geq 5$, $C$ is not hyperelliptic, and not trigonal, then by $\forall p\in C$, $\xi_{p}$ is not an asymptotic direction.
In fact, in [5] (see also [3, Theorem 2.2]) it is proven that for any $Q\in I_{2}(K_{C})$, we have
@@
and by [2, Theorem 6.1] we know that if $C$ is not hyperelliptic, and not trigonal of genus $g\geq 5$, for any $p\in C$, there exists a quadric $Q\in I_{2}(K_{C})$ such that $\mu_{2}(Q)(p)\neq 0$.
Assume that $C$ is trigonal not hyperelliptic and $g\geq 5$, then a basis for $I_{2}(K_{C})$ is given by the quadrics of rank at most 4 defined as in (11) where $L$ is the $g^{1}_{3}$. For such quadrics we have
@@
(see [5] for the first equality, and [2, Lemma 2.2] for the second one).
@@
Assume $C$ trigonal and $p$ not a ramification point of the $g^{1}_{3}$, then $h^{0}(K_{C}\otimes L^{\vee}(-p))=h^{0}(K_{C}\otimes L^{\vee})-1$, otherwise $|L(p)|$ is a $g^{2}_{4}$ and $C$ would be hyperelliptic. We claim that if $g\geq 6$, $h^{0}(K_{C}\otimes L^{\vee}(-2p))=h^{0}(K_{C}\otimes L^{\vee}(-p))-1$. In fact, otherwise $|L(2p)|$ would be a $g^{2}_{5}$, so it would give a map $C\to{\mathbb{P}}^{2}$ of degree 1, hence $g\leq 6$. So $g=6$ and $C$ would be a smooth plane quintic, hence not trigonal. So if we take $\tau_{1},\tau_{2}\not\in H^{0}(K_{C}\otimes L^{\vee}(-2p))\subset H^{0}(K_{C}\otimes L^{\vee})$, we get
@@
since $p$ is not a ramification point for $|L|$, nor for the pencil $\langle\tau_{1},\tau_{2}\rangle\subset H^{0}(K_{C}\otimes L^{\vee})$.
The case $Cliff(C)>1$ has already been considered in Theorem 5.5 where we proved, in particular, that there are no asymptotic directions of rank 1. So we assume $Cliff(C)=1$. Hence either $C$ is trigonal, or it is a smooth plane quintic. Assume first $C$ is trigonal.
@@
By the discussion following Remark 8.1, the deformations of rank one correspond to the points of ${\mathbb{P}}^{2}$, hence they are the intersections of two lines.
@@
Since the Schiffer variations are not asymptotic directions (see Lemma 8.2), we consider only rank one deformations corresponding to points $p\in{\mathbb{P}}^{2}\setminus C$.
@@
Since $L^{\otimes 2}=K_{C}$, we have $H^{0}(K_{C}\otimes L^{-2})=H^{0}({\mathcal{O}}_{C})$, and we choose $t=1\in H^{0}({\mathcal{O}}_{C})$. Choose $\omega=s_{1}s_{2}$, set $D=div(s_{1})$. Then the element $\zeta=[{\bar{\partial}}(\frac{\rho_{D}}{\omega})]$ is double-split of a rank one, by Proposition 2.3. Hence by Proposition 6.2, we have $\rho(Q)(\zeta\odot\zeta)=deg(L)=5,$ so $\zeta$ is not asymptotic.

### v2

@@ 5. Deformations of rank $d<Cliff(C)$
Let $C$ be a smooth algebraic curve of genus $g\geq 4$ and $\zeta\in H^{1}(T_{C})$ a deformation of rank $d>0$. Suppose one of the following assumptions is satisfied:
@@
Hence we have $h^{0}(K_{C}\otimes A^{\vee})\geq 1$. We claim that $h^{0}(A)\geq 2$. In fact, if $h^{0}(A)\leq 1$, we would have $h^{0}(K_{C}\otimes A^{\vee})\geq h^{0}(E)-1=g-d$ and by Riemann Roch, we would get $deg(A)\leq d<\frac{g-1}{2}$, a contradiction. So $h^{0}(A)\geq 2$, and if $h^{0}(K_{C}\otimes A^{\vee})\geq 2$, $A$ contributes to the Clifford index and we have:
@@
as the locus of deformations of rank $\leq d$. When $d=1$ it is the bicanonical curve.
@@
Let $C$ be a smooth curve of genus $g$, take a positive integer $d<Cliff(C)$ and an effective divisor $D$ of degree $d$. Then $K_{C}(-D)$ is projectively normal, so the multiplication map
@@
Notice that $h^{0}(D)=1$ since $Cliff(C)>d$. Moreover $K_{C}(-D)$ is very ample by Riemann Roch and by the assumption $Cliff(C)>d$. So by [19, Theorem 1], $K_{C}(-D)$ is projectively normal. ∎
@@
Let $C$ be a smooth curve of genus $g$, take a positive integer $d<Cliff(C)$, an effective divisor $D=\sum_{i=1}^{k}m_{i}p_{i}$ of degree $d$. Then any linear combination of all Schiffer variations $\xi^{n_{i}}_{p_{i}}$, with $n_{i}\leq m_{i}$ is not an asymptotic direction.
@@
We will prove the statement by induction on $d$. The case $d=1$ is already known. In fact, in [9] (see also [7, Theorem 2.2]) it is proven that for any $Q\in I_{2}(K_{C})$, we have
$II(Q)(\xi_{p}\odot\xi_{p})=-2\pi i\mu_{2}(Q)(p),$
and by [6, Theorem 6.1] we know that if $C$ is not hyperelliptic, and not trigonal of genus $g\geq 5$, for any $p\in C$, there exists a quadric $Q\in I_{2}(K_{C})$ such that $\mu_{2}(Q)(p)\neq 0$.
By induction assume that no linear combinations of all $\{\xi^{l_{i}}_{p_{i}}\ |\ l_{i}\leq m_{i},\ i<k,\ l_{k}\leq m_{k}-1\}_{i=1,...,k}$ are asymptotic.
@@
Let $C$ be a smooth curve of genus $g$, take a positive integer $d<Cliff(C)$. Then there are no asymptotic directions $\zeta\in H^{1}(T_{C})$ of rank $d$.
@@
    all tangent directions $\zeta$ of rank $d$, with $0<d<\lfloor\frac{g-1}{2}\rfloor$, are linear combinations of Schiffer variations supported on an effective divisor of degree $d$.
@@
    There are no asymptotic directions of rank $d$, with $0<d<\lfloor\frac{g-1}{2}\rfloor$.
@@
For the general curve of genus $g$ we characterise the image of the natural map ${\mathbb{P}}T_{C_{d}}\to{\mathbb{P}}(H^{1}(T_{C}))$ as the locus of deformations of rank at most $d$, for all $0<d<\lfloor\frac{g-1}{2}\rfloor$.
@@ 6. Double-split deformations
A nonzero deformation $\zeta\in H^{1}(T_{C})$ is said to be double-split if it is split with $\omega=s_{1}\tau$, $s_{1}\in H^{0}(L)$, $\tau\in H^{0}(K_{C}\otimes L^{\vee})$ and $\tau=s_{2}t$ with $s_{2}\in H^{0}(L)$ and $t\in H^{0}(K_{C}\otimes L^{-2})$.
@@
With the above notation, if $\zeta$ is double-split, and $Q=(s^{2}_{1}t)\odot(s_{2}^{2}t)-(s_{1}s_{2}t)\odot(s_{1}s_{2}t),$ we have
$II(Q)(\zeta\odot\zeta)=2\pi i(deg(L))\neq 0.$
@@
$II(Q)(\zeta\odot\zeta)=-2\pi i\sum_{p\in Supp(D)}Res_{p}(g_{1}dg_{2}-g_{3}dg_{4})=-2\pi i\sum_{p\in Supp(D)}Res_{p}(g_{1}dg_{2})=$
$=-2\pi i\sum_{p\in Supp(D)}Res_{p}(xd(\frac{1}{x}))=2\pi i\sum_{p\in Supp(D)}Res_{p}(\frac{dx}{x})=2\pi i(deg(D))=2\pi i(deg(L)).$
@@ 7. Rank $d=Cliff(C)$ deformations
Taking gobal sections in (1), if $\zeta$ has rank $d$, we get $h^{0}(E)=g-d+1$.
@@
Assume $C$ is not hyperelliptic. Recall that if $\zeta\in H^{1}(T_{C})$ is a non trivial deformation of rank $d=Cliff(C)<\frac{g-1}{2}$ and $E$ is not globally generated, then by Theorem 5.1, $\zeta$ is a linear combination of Schiffer variations supported on a degree $d$ effective divisor.
@@
Assume $\zeta\in H^{1}(T_{C})$ is an infinitesimal deformation of positive rank $d$ on a curve $C$ of Clifford index $d\neq\frac{g-1}{2}$ and $g\geq 5$. Assume moreover that $Cliff(C)=Cliff(L)=gon(C)-2$, where $L$ is a $g^{1}_{d+2}$ on $C$. Assume that $\zeta$ is not a linear combination of Schiffer variations supported on a degree $d$ effective divisor.
@@
By the assumptions and by Remark 7.1, we know that $E$ is globally generated. Take two generic points $p,q\in C$, then since $g\geq 5$, there exists a nontrivial holomorphic section $\sigma\in H^{0}(E(-p-q))$.
@@
where $E^{\prime}=E\otimes{\mathcal{O}}_{C}(-p-q)$, so $h^{0}(E^{\prime})\geq g-d-3>0$. By the theorem of Segre-Nagata and Ghione, ([24] p. 84), there exists a subline bundle $A$ of $E^{\prime}$ such that $deg(A)\geq\frac{g-5}{2}$. We claim that $h^{0}(A)\geq 1$. In fact, if $h^{0}(A)=0$, by Riemann Roch we would have $h^{0}(K_{C}\otimes A^{\vee})=g-1-deg(A)\leq\frac{g+3}{2}$. So, since $p,q$ are generic points, $h^{0}(K_{C}\otimes A^{\vee}(-2p-2q))\leq\frac{g-5}{2}$. On the other hand, up to saturation of $A$, so that the quotient $E^{\prime}/A$ is torsion free, we have the exact sequence
@@
We have $h^{0}(A(p+q))\geq 1$ and $deg(A(p+q))\geq\frac{g-1}{2}$. If $h^{0}(A(p+q))=1$, then $h^{0}(K_{C}\otimes A^{\vee}(-p-q))=g-2-deg(A)\leq\frac{g+1}{2}$, but $h^{0}(K_{C}\otimes A^{\vee}(-p-q))\geq h^{0}(E)-1=g-d$, so we get $d\geq\frac{g-1}{2}$, a contradiction. So $h^{0}(A(p+q))\geq 2$ and $h^{0}(K_{C}\otimes A^{\vee}(-p-q))\geq 2$, since $E$ is globally generated, hence also $K_{C}\otimes A^{\vee}(-p-q)$ is globally generated and non trivial.
So $A(p+q)$ contributes to the Clifford index and we have:
@@
hence $d\leq Cliff(A(p+q))\leq d$, and either $A(p+q)=L$, or $A(p+q)=K_{C}\otimes L^{\vee}$, where $L$ is any line bundle as in the statement. But $A(p+q)\neq L$, since $p,q$ are general and $A$ is effective. So $A(p+q)=K_{C}\otimes L^{\vee}$, and the vertical exact sequence induces an exact sequence on global sections. So it corresponds to a class of an extension $\epsilon\in H^{1}(K_{C}\otimes L^{-2})$ that induces the zero map $H^{0}(L)\to H^{1}(K_{C}\otimes L^{\vee})$. The map $H^{1}(K_{C}\otimes L^{-2})\cong H^{0}(L)^{\vee}\rightarrow Hom(H^{0}(L),H^{1}(K_{C}\otimes L^{\vee}))\cong H^{0}(L)^{\vee}\otimes H^{0}(L)^{\vee}$, $v\mapsto\cup v$, is the dual of the multiplication map $H^{0}(L)\otimes H^{0}(L)\rightarrow H^{0}(L^{\otimes 2})$ and $\epsilon$ is in the kernel of this map. If $h^{0}(L^{\otimes 2})=3$, by the base point free pencil trick, the multiplication map $Sym^{2}H^{0}(L)\to H^{0}(L^{\otimes 2})$ is surjective, so $\epsilon$ must be zero. Hence $E=(K_{C}\otimes L^{\vee})\oplus L$ and $\zeta$ is split.
@@
Assume now $h^{0}(L^{\otimes 3})=4$, then the multiplication map
@@ 8. Rank one deformations
Assume $C$ is not hyperelliptic, $\zeta\in H^{1}(T_{C})$ a rank one deformation. If $E$ is not globally generated, then $\zeta=\xi_{p}$ is a Schiffer at a point $p$. If $C$ is non trigonal, then $\forall p\in C$, $\xi_{p}$ is not an asymptotic direction. If $C$ is trigonal and $g\geq 6$, then $\xi_{p}$ is an asymptotic direction if and only if $p$ is a ramification point of the $g^{1}_{3}$.
@@
Assume $E$ is not globally generated, we conclude by Theorem 5.1 that $\zeta=\xi_{p}$ is a Schiffer variation at a point $p$.
If $g\geq 5$, $C$ is not hyperelliptic, and not trigonal, then for any point $p\in C$, $\xi_{p}$ is not an asymptotic direction. In fact, in [9] (see also [7, Theorem 2.2]) it is proven that for any $Q\in I_{2}(K_{C})$, we have
@@
and by [6, Theorem 6.1] we know that if $C$ is not hyperelliptic, and not trigonal of genus $g\geq 5$, for any $p\in C$, there exists a quadric $Q\in I_{2}(K_{C})$ such that $\mu_{2}(Q)(p)\neq 0$.
Assume that $C$ is trigonal not hyperelliptic and $g\geq 5$, then a basis for $I_{2}(K_{C})$ is given by the quadrics of rank at most 4 defined as in (11) where $L$ is the $g^{1}_{3}$ (see Remark 6.1).
For such quadrics we have
@@
(see [9] for the first equality, and [6, Lemma 2.2] for the second one).
@@
Assume $C$ trigonal and $p$ not a ramification point of the $g^{1}_{3}$, then $h^{0}(K_{C}\otimes L^{\vee}(-p))=h^{0}(K_{C}\otimes L^{\vee})-1$, otherwise $|L(p)|$ is a $g^{2}_{4}$ and $C$ would be hyperelliptic. We claim that if $g\geq 6$, $h^{0}(K_{C}\otimes L^{\vee}(-2p))=h^{0}(K_{C}\otimes L^{\vee}(-p))-1$. In fact, otherwise $|L(2p)|$ would be a $g^{2}_{5}$, so it would give a map $C\to{\mathbb{P}}^{2}$ of degree 1, hence $g\leq 6$. So $g=6$ and $C$ would be a smooth plane quintic, hence not trigonal. So if we take a pencil $\langle\tau_{1},\tau_{2}\rangle\subset H^{0}(K_{C}\otimes L^{\vee})$ such that $\langle\tau_{1},\tau_{2}\rangle\cap H^{0}(K_{C}\otimes L^{\vee}(-2p))=\{0\}$, we get
@@
since $p$ is not a ramification point for $|L|$, nor for the pencil $|\langle\tau_{1},\tau_{2}\rangle|$.
@@
By the discussion following Remark 8.1, the deformations of rank one correspond to the points of ${\mathbb{P}}^{2}$, hence they are the intersections of two lines. Since the Schiffer variations are not asymptotic directions (see Lemma 8.2), we consider only rank one deformations corresponding to points $p\in{\mathbb{P}}^{2}\setminus C$.
@@
Since $L^{\otimes 2}=K_{C}$, we have $H^{0}(K_{C}\otimes L^{-2})=H^{0}({\mathcal{O}}_{C})$, and we choose $t=1\in H^{0}({\mathcal{O}}_{C})$. Choose $\omega=s_{1}s_{2}$, set $D=div(s_{1})$. Then the element $\zeta=[{\bar{\partial}}(\frac{\rho_{D}}{\omega})]$ is double-split of rank one, by Proposition 2.3. Hence by Proposition 6.3, we have $\rho(Q)(\zeta\odot\zeta)=(2\pi i)deg(L)=10\pi i,$ so $\zeta$ is not asymptotic.